If the perimeter of a rectangle is 10a-6 metres and the width is 2a metres, what do I do to find out the length of this rectangle?
Answer by waynest(281)

p = 10a - 6
width = 2a
length = L
p = 2L + 2w
2(2a) + 2L = 10a - 6
4a + 2L = 10a - 6
4a - 4a + 2L = 10a - 4a - 6
2L = 6a - 6
2L/2 = (6a 6)/2
L = 3a - 3
check:
2(2a) + 2(3a - 3) = 120a - 6
4a + 6a - 6 = 10a - 6
10a - 6 = 10a - 6